what is the next number and how did you arrive at it? 7,16,8,27,9, is the answer 27 or 10 or 40 or 37
Answer by macston(5194)

Odd terms are counting upward 7,8,9,...
Even terms are the product of (previous term+1) and (previous term-5)
7...(7+1)*(7-5)=
16
8...(8+1)*(8-5)=
27
9...(9+1)*(9-5)=
40
ANSWER: The next number would be 40.